You are doing and experiment on inbreeding in Phlox plants with a single-locus polymorphism in flower color. WW genotypes have white flowers, PP genotypes have purple flowers, and WP genotypes have magenta flowers. You construct an artificial population in the greenhouse consisting of 50 white-flowered phenotypes and 50 purple-flowered phenotypes. You then cross all individuals randomly and grow out 100 F1 seeds. What are the genotype frequencies among the F1 offspring if you assume Hardy Weinberg equilibrium?
Answer #1

Answer:

I am doing and experiment on inbreeding in Phlox plants with a single-locus polymorphism in flower color.

WW genotypes have white flowers, PP genotypes have purple flowers, and WP genotypes have magenta flowers.

I construct an artificial population in the greenhouse consisting of 50 white-flowered (WW) phenotypes and 50 purple-flowered (PP) phenotypes.

So, the frequency of WW genotypes = number of WW genotypes / total number of all genotypes = 50/100 = 0.5 and the frequency of PP genotypes = number of PP genotypes / total number of all genotypes = 50/100 = 0.5

Crossing is done among all individuals randomly and grows out 100 F1 seeds. The genotype frequencies among the F1 offspring if I assume Hardy Weinberg equilibrium is—

WW (0.5)

PP (0.5)

WW (0.5)

WW×WW (0.25)

PP×WW (0.25)

PP (0.5)

WW×PP (0.25)

PP×PP (0.25)

Mating of genotype

Total frequency

Frequency of WW

Frequency of WP

Frequency of PP

WW×WW

0.25

0.25

WW×PP

0.25

0.25

PP×WW

0.25

0.25

PP×PP

0.25

0.25

Total®

1.0

0.25

0.5

0.25

So, the genotype frequencies among the F1 offspring if I assume Hardy Weinberg equilibrium is— WW = 0.25, WP = 0.5 and PP = 0.25